A rectangular painting is W=117cm wide and H=76.2cm high. At what speed,v,must the painting move parallel to its width if it is to appear to be square? Please show steps.
Answer #1

For the painting to appear square, the width shoulf apparently appear to be = H=76.2 cm . This is because only that dimension apparently contracts parallel to which the painting is moving.

Let the required speed be v

H=W*sqrt(1 - v²/c²)

76.2/117 = sqrt(1-v²/9*10^16)

Solving for v,we get

v=2.28*10^8 m/s which is approximately 0.76c